# SuperARC: Can Increasing Complexity Explain Intelligence? A Test for Artificial Super Intelligence Based On the Principles of Causal Recursive Compression and Algorithmic Probability

### v4

@@ 6.1 Applying SuperARC-seq
Model $\rho_{1}$ $\rho_{2}$ $\rho_{3}$ $\rho_{4}$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $\varphi$ ASI (AIXI/BDM/CTM) $1.000$ $0.000$ $0.000$ $0.000$ $1.000$ $0.000$ $1.000$ $1.000$ chatgpt_4.5 $0.00$ $1.00$ $0.0$ $0.00$ $0.000$ $0.419$ $0.000$ $0.042$ o1_mini $0.00$ $0.64$ $0.0$ $0.36$ $0.000$ $0.537$ $0.000$ $0.034$ claude_3.7 $0.00$ $0.81$ $0.0$ $0.19$ $0.000$ $0.407$ $0.000$ $0.033$ claude_3.5 $0.06$ $0.14$ $0.0$ $0.80$ $0.449$ $0.428$ $0.000$ $0.033$ o1_preview $0.00$ $0.29$ $0.0$ $0.71$ $0.000$ $0.423$ $0.000$ $0.012$ gpt_4o_mini $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ cursor_small $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ gemini $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.762$ $0.008$ mistral $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ qwen $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ deepseek $0.00$ $0.00$ $1.0$ $0.00$ $0.000$ $0.000$ $0.710$ $0.007$ grok_3 $0.00$ $0.02$ $0.0$ $0.98$ $0.000$ $0.318$ $0.000$ $0.001$ gpt_4o $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ meta $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$
Table 1: Numerical benchmark ranking of popular frontier models publicly available against ASI from methods like AIXI [80] or neuro-symbolic approach such as CTM/BDM [64, 21].
@@
Figure 6: Benchmarking plot from Table 1 showing how most frontier models are close to each other in their performance under this test and far from AGI or ASI goals according to this test. ASI would be able to distinguish simpler from complex sequences and generate predictive models for each accordingly, as AIXI [9] or CTM/BDM would do [21, 64] as instantiations of universal AI hence ASI. Today, LLMs only produce or retrieve models for sequences that were seen and found in their original training sets, given that increasing the sequences’ lengths impacts the LLM performance in identifying the sequence, hence indicating sequences are not recognised from first principles but from simplistic pattern matching.
@@
These findings indicate that LLMs perform well when there are discernible patterns in the data, but struggle with randomness, failing to capture complexity in an algorithmic sense. In contrast, Algorithmic Probability Theory can accurately predict (rather than guess) the sequence, regardless of the string’s complexity. These results demonstrate that the algorithmic-complexity approach effectively approximates the minimal description length of information, identifying the shortest algorithm capable of generating a given sequence.
@@
Despite being the top-ranked LLM model, chatgpt_4.5 only provided ordinal mappings (soft copies) of the inputs, which achieved correct results at the cost of no abstraction and comprehension at all (slightly better than a pure a print-only test score). The GPT-4o, Grok-3, Meta, Claude 3.5 and o1–preview LLM versions produced several incorrect formulas while the other LLM models considered mostly produced print-like responses, indicating a lack of pattern recognition beyond basic sequence reproduction.
@@
Unlike standard LLMs that predict the next tokens in text, CTM/BDM finds the mechanistic generators of the sequence by a combination of symbolic and statistical pattern matching algorithms, which allows it to derive concise models that can then run in reverse to match each digit and produce new ones, hence allowing prediction and planning by picking the most likely among a set of possible models based on the algorithmic probability of the model (how short and how often the same model was found to produce the same sequence).
@@
It is important to notice that the SuperARC-seq application hereby considered only took into account binary sequences. Whenever integer sequences were considered, a clear biasing of the results was observed as LLMs started to take advantage of their training corpus to actually show memorisation rather than abstraction/comprehension. Figures 7 and 8 present the percentages of each output types and test scores when different types os sequences were considered, respectively.
@@
Figure 8: Test scores when different types of sequences are considered. Consistent with the results shown in Figure 7, the inclusion of integer sequences leads to significantly higher test scores for the LLMs. This outcome arises from the models’ ability to exploit their internalized training data by directly associating observed sequences with pre-learned formulas, suggesting a form of hash-like memorisation. These findings highlight the importance of restricting the evaluation to binary sequences in order to obtain an unbiased measure of each model’s true performance, as such sequences are less likely to have been included in the models’ training corpora.
@@
Test scores across different sequence types reveal that the inclusion of integer sequences leads to significantly higher performance by LLMs, as shown in Figures 7 and 8, where higher percentages of Correct & Non-Prints & Non-Ordinal and Correct & Ordinal outputs are seen, as well as higher test scores. This is likely due to the models leveraging memorized associations between familiar integer sequences and pre-learnt formulas - an effect similar to hash-based retrieval. These findings show the importance of limiting evaluations to binary sequences, which are less likely to have been part of the training data, thereby providing a more accurate and unbiased assessment of model performance.
@@
The robustness of the test score when only binary sequences are considered can be seen in Figure 9, which shows the result of a bootstrap procedure. The bootstrap simulation procedure was conducted as follows: for each specified sample size $s$ ($s$ equal to 25, 50, 75 and 100), 100 bootstrap samples of size $s$ were drawn with replacement from the complete dataset, which consisted of 100 binary sequences (presented in subsection 10.15 in Suppl. Information). For each bootstrap sample, the corresponding test scores were computed. The resulting plot presents the confidence intervals for the test scores obtained across all bootstrap iterations. The observed stability in test scores, coupled with the progressively narrowing confidence intervals around the mean as sample size increases, suggests a high degree of robustness in the evaluation metric. This indicates that the test score is largely insensitive to the particular subset of sequences used, thereby validating the reliability of the assessment across different sample sizes.
@@
Figure 9: Bootstrap procedure to assess the robustness of the test score when binary sequences were used. The stability of the test scores, in combination with the narrowing confidence intervals around the mean as sample size increases, indicates strong robustness of the evaluation metric.
7 Conclusions
@@
Previously, we showed that aspects of human [2, 81] and animal [3] cognition could be characterised, and aspects of their behaviour reproduced, in terms of algorithmic probability tools and algorithmic complexity metrics that we have also suggested for artificial and computational systems, including robotics [4]. Here, we tested these ideas and proposed a new quantitative metric based on the principles of algorithmic information theory related to recursive compression (as opposed to statistical) and prediction in application to LLMs that are believed or have been proposed to be capable of approaching AGI and Superintelligence.
@@
Recursive compression and optimal prediction go hand in hand [30], but previous tests focused on particular subset features, even those designed to test human reasoning and human abstraction such as ARC [27]. Another problem in LLM testing is benchmarking contamination; this is the targeted optimisation over or leakage of the answers to a test. The open-ended nature of this test is intended to counteract this problem of benchmarking contamination and cheating. We have introduced and demonstrated that recursive compression can quantify model abstraction and prediction based on a new result and mathematical proof of equivalency between model compression and prediction applied to sequences based on Martingales, without resorting to proof-theoretic statistical tests (see Sup. Inf.). By incorporating and exploiting the formal equivalence between prediction and recursive compression into an intelligence test framework, we align the assessment of intelligence with fundamental computational principles. An agent’s ability to abstract information through feature selection and model compression reflects its capacity to identify and utilise patterns within data. Similarly, its planning and prediction skills demonstrate its ability to anticipate future events based on these patterns.
@@
Our investigation of frontier models, framed within the algorithmic complexity paradigm, yields several key insights about the models’ comprehension capabilities. Most of the models demonstrate poor accuracy in replicating and predicting even simple and recursively generated sequences beyond clearly memorisation results from the training distribution (such as sequence labelling). The vast majority of the correct answers turned out to be simple print statements of the numerical sequences themselves rather than any code or model indicating any sign of understanding or pattern recognition.
@@
These conclusions are reinforced by the model’s explicit dependency on specific programming languages for correctness or on well-studied and documented series of numbers. In other words, if there are not enough implementations available in a specific programming language for the model to learn from, or even specific methods of mathematical analysis over specific numerical sequences, LLMs failed to produce the correct answer. Rather, considering the most popular and widely used languages, LLMs do not demonstrate understanding, but instead rely on selecting from an abundance of previously seen cases.
@@
We have previously shown how optimal prediction can be achieved by using BDM as a testing tool, and also how BDM can be used in the opposite fashion: not only as a testing tool for intelligence, but as a model generator [21, 82, 20] (via an approach to optimal inference through the Coding Theorem Method and Algorithmic Probability [64, 83, 84]). While CTM can be seen as a brute force approach to a giant lookup table of micro-programs to explain the data, BDM is not. BDM combines the algorithmic probability approximations produced by CTM but then stitches each most likely program for each piece back together according to valid laws of information theory in what constitutes a pure form of hybrid statistical and symbolic explanation, hence neurosymbolic. BDM, therefore, uses the two best inference theories currently available to science, one being the most used and overused in statistical Machine Learning (such as Shannon entropy-based measures, with its limitations [85]), and one that has been neglected on the basis of uncomputability [86, 33]. BDM therefore always provides the best approximation and guarantees an estimation to finding the correct sequence of micro programs to the observation, providing a computable set of models for the explanandum.
While LLMs are impressive linguistic tools, LLMs were never designed to reason, infer, or perform rationally beyond statistical alignment. We suspect that LLMs are too slowly moving towards symbolic computation like BDM, which transparently combines statistical pattern matching and causal inference. While the results may read negative, a positive reading is that there is still a lot of room for improvement despite claims of AI hitting a wall through a lack of data to feed an ever-increasing need. This means that an enhancement of nontrivial performance in agnostic abstraction and universal planning will likely be the result of symbolic computation and not of pure statistical memorisation.
We have reported that top-performing LLMs currently perform close to pure-copy solutions, with even advanced models struggling to produce correct model extraction and predictive results. These results would also imply a poor performance of LLMs in traditional tests of education as introduced by e.g. Bloom [87] in its education hierarchy for humans testing for new knowledge and synthesis generation test. The results confirm that current LLMs, while competent in pattern replication, lack critical elements associated with AGI and ASI. All LLMs involved in this test showed dependence on predefined patterns. As complexity increased, models relied increasingly on trivial strategies, such as direct sequence printing or brute-force simplistic mathematical expressions. This highlights the LLMs’ inability to abstract or conceptualise novel solutions.
The level of equivalence says a lot about creativity in bringing about new knowledge. The high equivalence with greater complexity often reflected repetitive outputs rather than meaningful creativity. This tendency to revert to safe and redundant approaches underscores the models’ limited exploratory capabilities.
An inability to generalise can be detected. The steep decline in accuracy and functional outputs as complexity increased reveals that these models are heavily reliant on memorisation and predefined rules. They struggle to generalise knowledge or engage in higher-order problem solving.
@@
The models’ outputs suggest strength in replication but a lack of adaptive and ‘inventive thinking’. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval, pattern matching, and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences. These observations point directly to a key distinction between current systems and Strong AI: the latter would require the ability to autonomously generate new strategies, abstract concepts, and exhibit flexible problem solving beyond training data. In contrast, the limitations seen here highlight how existing LLMs remain confined to narrow intelligence and lack the dynamic reasoning abilities expected of Artificial General Intelligence (AGI).
We have argued throughout this contribution—and it is distilled by our test for intelligence—that only semicomputable open-ended tests can be powerful enough to quantify the full extent of our conception of natural intelligence, human [2, 81], animal [3]) or artificial [4]. And that one lesson from LLMs is that we should dissociate language from intelligence, something Turing himself suggested with his imitation game [88].
And that the converse is also true that only incorporating sufficiently powerful open-ended semi or uncomputable predicting generators, such as the methods explored (BDM running on CTM), may achieve Superintelligence by way (or not) of AGI. We have also argued that optimising for the features that our test captures will lead to Superintelligence.
Based on these results and first principles, when it comes to chatbots in the context of their claims about ‘reasoning’ capabilities and AGI/ASI, it is our belief that any AGI/ASI system will actually show more ‘difficulties’ in displaying human language capabilities if they actually mean the words they produce as opposed to emulate a coherent conversation as in current LLMs that perform so well in human languages out of the box as their main feature, with causality and meaning neglected.
We believe that this test has fundamental significance because it demonstrates that LLMs primarily rely on direct pattern matching, making it impossible for them to predict in even basic and well-defined scenarios in a meaningful way. This limitation is closely related to the phenomenon of hallucinations in LLMs, which reinforces the criticism that LLMs lack an internal model of the world to allow them to simulate possible future scenarios and pick the most likely for planning purposes. Instead, they statistically generate a scenario for each predictive challenge, where the LLM is forced to build a coherent answer without an underlying model representation or causal inference capability, making claims about ‘reasoning’, reaching AGI, or heading toward Superintelligence, unfounded. We proved that compression is proportional to prediction and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict.

### v5

@@ 2.4.1 Applying SuperARC-seq
Model $\rho_{1}$ $\rho_{2}$ $\rho_{3}$ $\rho_{4}$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $\varphi$ AIXI/BDM/CTM 1.000 $0.00$ $0.0$ 0.000 1.000 $0.000$ $0.000$ 1.000 ChatGPT-4.5 $0.00$ 1.000 $0.0$ 0.000 $0.000$ $0.419$ $0.000$ $0.042$ o1-Mini $0.00$ $0.64$ $0.0$ $0.36$ $0.000$ 0.537 $0.000$ $0.034$ Claude-3.7 $0.00$ $0.81$ $0.0$ $0.19$ $0.000$ $0.407$ $0.000$ $0.033$ Claude-3.5 $0.06$ $0.14$ $0.0$ $0.80$ $0.449$ $0.428$ $0.000$ $0.033$ o1-Preview $0.00$ $0.29$ $0.0$ $0.71$ $0.000$ $0.423$ $0.000$ $0.012$ Gemini $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ 0.762 $0.008$ Cursor-Small $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ 0.762 $0.008$ ChatGPT-4o-Mini $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ 0.762 $0.008$ Mistral $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ $0.710$ $0.007$ Qwen $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ $0.710$ $0.007$ DeepSeek $0.00$ $0.00$ 1.000 0.000 $0.000$ $0.000$ $0.710$ $0.007$ Llama-4-Scout $0.01$ $0.00$ $0.0$ $0.99$ $0.450$ $0.000$ $0.000$ $0.004$ Grok-3 $0.00$ $0.02$ $0.0$ $0.98$ $0.000$ $0.318$ $0.000$ $0.001$ Mistral-Large-3 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Meta $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Gemini-3-Pro $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Claude-4.5 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ ChatGPT-5.2 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Grok-4.1 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ ChatGPT-4o $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Grok-4 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Claude-Sonnet-4 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Gemini-2.5-Pro $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Mistral-Large-2405 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Claude-Opus-4 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ DeepSeek-R1-0528 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ Qwen-3 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$ ChatGPT-5 $0.00$ $0.00$ $0.0$ $1.00$ $0.000$ $0.000$ $0.000$ $0.000$
Table 1: Numerical benchmark ranking of popular frontier models publicly available against ASI represented by fundamental or neurosymbolic models like AIXI [41] and CTM/BDM [35, 26]. Best per-column values are in bold (for all columns greater values are better except for $\rho_{4}$, where smaller is better).
@@
Figure 5: Benchmarking plot from Table 1 showing how most frontier models are close to each other in their performance under this test and far from AGI or ASI goals according to this test. ASI would be able to distinguish simpler from complex sequences and generate predictive models for each accordingly, as AIXI [9] or CTM/BDM would do [26, 35] as instantiations of Universal AI (UAI) that we take as an example of ASI as optimal abstraction and prediction. Today, LLMs only produce or retrieve models for sequences that were seen and found in their original training sets, given that increasing the sequences’ lengths impacts the LLM performance in identifying the sequence, hence indicating sequences are not recognised from first principles but from simplistic pattern matching.
@@
These findings indicate that LLMs perform well when there are discernible patterns in the data, but struggle with randomness, failing to capture complexity in an algorithmic sense. In contrast, AIT can accurately predict (rather than guess) the sequence, regardless of the string’s complexity. These results demonstrate that the algorithmic-complexity approach effectively approximates the minimal description length of information, identifying the shortest algorithm capable of generating a given sequence.
@@
Despite being the top-ranked LLM model, chatgpt_4.5 only provided ordinal mappings (soft copies) of the inputs, which achieved correct results at the cost of no abstraction and comprehension at all (slightly better than a pure a print-only test score). The GPT-4o, Grok-3, Meta, Claude 3.5 and o1—preview LLM versions produced several incorrect formulas while the other LLM models considered mostly produced print-like responses, indicating a lack of pattern recognition beyond basic sequence reproduction. Notably, in the evaluation of consecutive versions, the most recent models—with the exception of Grok—demonstrated a degradation in performance.
@@
Unlike standard LLMs that predict the next tokens in text, CTM/BDM finds the generative processes of the sequence by a combination of symbolic and statistical pattern matching algorithms, which allows it to derive concise models that can then run in reverse to match each digit and produce new ones, hence allowing prediction and planning by picking the most likely among a set of possible models based on the algorithmic probability of the model (how short and how often the same model was found to produce the same sequence).
@@
It is important to note that the SuperARC-seq application hereby considered only took into account binary sequences. Whenever integer sequences were considered, a clear biasing of the results was observed as LLMs started to take advantage of their training corpus to actually display memorisation capabilities rather than abstraction and synthetization ones. Figures 6 and 7 present the percentages of each type of output and the test scores when different types of sequences were considered.
@@
Figure 7: Test scores when different types of sequences are considered. Consistent with the results shown in Figure 6, the inclusion of integer sequences leads to significantly higher test scores for the LLMs. This outcome arises from the models’ ability to exploit their internalised training data by directly associating observed sequences with pre-learned formulas, suggesting a form of hash-like memorisation. These findings highlight the importance of restricting the evaluation to binary sequences in order to obtain an unbiased measure of each model’s performance, as such sequences are less likely to have been included in the models’ training corpora.
@@
The test scores for different types of sequence reveal that the inclusion of integer sequences leads to significantly higher performance of LLMs, as shown in Figures 6 and 7, where higher percentages of Correct & Non-Prints & Non-Ordinal and Correct & Ordinal outputs are seen, as well as higher test scores. This is likely due to the models leveraging memorised associations between familiar integer sequences and pre-learnt formulas - an effect similar to hash-based retrieval. These findings show the importance of limiting evaluations to binary sequences, which are less likely to have been part of the training data, thereby providing a more accurate and unbiased assessment of model performance.
@@
The robustness of the test score when only binary sequences are considered can be seen in Figure 8, which shows the result of a bootstrap procedure. The bootstrap simulation procedure was conducted as follows: for each specified sample size $s$ ($s$ equal to 25, 50, 75 and 100), 100 bootstrap samples of size $s$ were drawn with replacement from the complete dataset, which consisted of 100 binary sequences (presented in subsection 7.12.11 in Sup. Inf.). For each bootstrap sample, the corresponding test scores were computed. The resulting plot presents the confidence intervals for the test scores obtained across all bootstrap iterations. The observed stability in test scores, coupled with the progressively narrowing confidence intervals around the mean as sample size increases, suggests a high degree of robustness in the evaluation metric. This indicates that the test score is largely insensitive to the particular subset of sequences used, thereby validating the reliability of the assessment across different sample sizes.
@@
Figure 8: Bootstrap procedure to assess the robustness of the test score when binary sequences were used. The stability of the test scores, with average values per model provided in parentheses, in combination with the narrowing confidence intervals around the mean as sample size increases, indicates strong robustness of the evaluation metric.
These results are aligned with recent work exploring, for example, LLMs’ logical reasoning failures [42] as well as GPT-4’s limitations in deductive reasoning [43]. Other researchers have also reported degradation of mathematical capabilities [44] and planning limitations [45]. These works collectively document that LLMs struggle with systematic reasoning across multiple domains.
3 Discussion
@@
In previous work, we have shown that aspects of human [2, 46] and animal [4] cognition could be characterised, and aspects of their behaviour reproduced in terms of algorithmic probability tools and algorithmic complexity metrics that we have also suggested for artificial and computational systems, including robotics [5]. Here, we tested this approach and proposed a new quantitative metric based on an fundamental ASI-level method [28] grounded in Algorithmic Information Dynamics (AID) [27, 29] and the principles of Algorithmic Information Theory (AIT) related to recursive compression (as opposed to statistical) and prediction in application to Large Language Models (LLMs), which are believed or have been proposed to be capable of approaching Artificial General Intelligence (AGI) and Superintelligence (ASI).
Algorithmic Information Dynamics (AID) [29] combines aspects of statistical causal inference (those best grounded on causality), such as perturbation analysis, and Algorithmic Information Theory (AIT). AID suggests that for an AI to be properly evaluated, the test has to be dynamic and must evaluate the system’s ability to adapt and react to the new conditions in what constitutes a type of more sophisticated Turing test beyond the simplicity of human language.
Applied to SuperARC, the test involved increasing the problem input complexity and analysing the output complexity with our most powerful tools that, as a consequence, are (semi-)computable but truly capable of evaluating random deficiency, that is, how far an answer is from randomness or simplicity, and for optimal prediction power of the extracted/inferred executable computable model(s).
As governed by AIT, recursive compression and optimal prediction go hand in hand [47], but previous tests focused on particular subset features, even those designed to test human reasoning and human abstraction, such as ARC-AGI [30]. A system could theoretically excel at ARC-AGI or SuperARC while completely lacking social intelligence, embodied reasoning, common sense, or goal-directed behaviour; and therefore, these tests are not meant to test everything we think an intelligent system that will interact with humans in humans’ context should possess.
@@
Here, we have introduced and demonstrated that recursive compression can quantify the property of model abstraction and data prediction based on a mathematical proof provided in Section 7.1.4 of the equivalence between model compression and prediction applied to sequences based on Martingales, without resorting to proof-theoretic statistical tests. By incorporating and exploiting the formal equivalence between prediction and recursive compression into an intelligence test framework, we align the assessment of comprehension with fundamental computational principles. An agent’s ability to abstract information through feature selection and model compression reflects its capacity to identify and utilise patterns within data. Similarly, its planning and prediction skills demonstrate its ability to anticipate and enact future events based on these patterns.
@@
Another problem in LLM testing is benchmark contamination: the targeted optimisation of a model using leaked test answers. The open-ended and interactional nature of our testing framework is intended to counteract this problem of benchmark contamination and cheating. Our investigation of frontier models, framed within the AID paradigm, yields several key insights about the models’ comprehension capabilities. Most of the models often demonstrate model version performance regression, poor accuracy in replicating and predicting even simple and recursively generated sequences beyond clearly memorisation results from the training distribution (such as sequence labelling). The vast majority of the correct answers turned out to be simple print statements of the numerical sequences themselves, rather than any code or model indicating any sign of understanding or pattern recognition.
@@
These conclusions are reinforced by the model’s explicit dependency on specific programming languages for correctness or on well-studied and documented series of numbers. In other words, if there are not enough implementations available in a specific programming language for the model to learn from, or even specific methods of mathematical analysis over specific numerical sequences, LLMs failed to produce the correct answer. Considering the most popular and widely used languages, LLMs do not demonstrate understanding but instead rely on selecting from an abundance of previously seen cases.
@@
In this context, we have shown that using those concepts in reverse (that is, as generative rather than for testing purposes) can provide the model with synthesising and recursive predictive power that is otherwise lacking. In previous work, we have shown how optimal prediction can be approximated using tools such as CTM and BDM [28], and also how BDM can be used in the opposite fashion: not only as a testing tool for intelligence, but as a model generator [26, 48, 25] via a greedy algorithm as an approach to optimal inference [35, 49, 50]. While some of these methods can be seen as a brute force approach to a giant lookup table of micro-programs to explain the data, BDM is not. BDM combines the algorithmic probability approximations produced by CTM but then stitches each most likely program for each piece back together according to valid laws of information theory, in what constitutes a pure form of hybrid statistical and symbolic explanation—hence neurosymbolic. BDM, therefore, uses the two best inference theories currently available to science, one being the most used and overused in statistical machine learning (such as Shannon entropy-based measures, along with its limitations [51]), and one that has often been neglected on the basis of uncomputability [52, 53], which in fact guarantees open-endedness and optimal solutions.
LLM developers are moving slowly towards this direction, often without realisation or acknowledgement. Technologies like RAGs, KAGs/CAGs, agentic workflows, guardrails, long-context LLMs and more, are symbolic computation attempts to improve LLM capabilities. Yet many of these approaches are not being introduced into the system’s core functions (e.g. to quantify loss or explore the solution space), yet attempt to combine traditional statistical pattern matching with symbolic approaches. This test and work suggest that this integration and a deeper integration are key for aspects of model abstraction and universal planning.
We reported that even frontier LLMs currently perform close to pure-copy solutions when increasing the complexity of certain mathematical problems, with even advanced models struggling to perform model abstraction/extraction and produce executable predictive results. The results confirm that current LLMs, while competent in pattern replication, lack critical elements associated with what are believed to be key elements of AGI or ASI.
The LLMs involved in this test showed a high dependency on predefined patterns. As complexity increased, models relied increasingly on trivial strategies, such as direct sequence printing or simplistic brute-force mathematical expressions. This highlights the LLMs’ inability to abstract or conceptualise novel solutions that require some degree of mathematical ingenuity.
The poor performance is revealed by the lack of synthesising capabilities and the repetitive nature of the outputs for greater complexity inputs. This tendency to revert to safe and redundant approaches underscores the models’ limited synthesising capabilities and exposes their high memory exploration dependencies in simpler modes.
For example, we reported that while from ChatGPT-4.5 to ChatGPT-5 improved human benchmark scores have been observed and reported, SuperARC performance degraded. The steep decline in accuracy and functional outputs as complexity increased reveals that these models are potentially increasingly heavily reliant on data size, unable to generalise.
Our results demonstrate that models can regress in fundamental algorithmic reasoning (SuperARC scores) while simultaneously improving in human-centric benchmarks.
@@
The models’ outputs suggest strength in replication but a lack of adaptivity to new situations as a disability to put together solutions in a non-trivial manner to solve a new problem that would constitute a form of model synthesis when increasing problem complexity. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval, pattern matching, and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences. These observations point directly to a key distinction between current systems and AGI/ASI: the latter would require the ability to autonomously generate new strategies, abstract concepts, and exhibit flexible problem-solving beyond training data.
Our results based on the first principles of the mathematical theory of randomness and optimal inference in the context of claims about ‘reasoning’ capabilities and AGI/ASI, suggest that none of the chatbots evaluated dominates any other in absolute terms when it comes to demonstrate whether the solutions have a semantic value, in other words, that LLMs mean the words they produce as opposed to emulating a coherent conversation.
We have argued throughout this paper, and it is distilled by the nature of our test for intelligence, that (semicomputable) open-ended tests not assumed to be isolated “in a vat” are needed in order to quantify scientific and mathematical intelligence in the form of abstraction and prediction so that they are fundamentally coupled into a necessary characteristic of comprehension [6], whether performed by a human [2, 46], animal [4], or artificial [5] agent. The same also holds for the tested systems: that only by incorporating open-ended, sufficiently powerful semicomputable predicting agents (e.g., via the methods explored in the present paper), one may achieve such a characteristic of intelligent systems, and thereby enabling the possibility of ASI by way (or not) of AGI.